Applying Bolzano-Weierstrass theorem show that the set
S={1+1/ n ∨ n∈N}∪{−1−1/ n ∨ n∈N}
must have a limit.
1
Expert's answer
2020-06-22T16:37:34-0400

{1+1/ n ∨ n∈N}

x1=1+1/1=2

x2=1+1/2=1.5

1.5-2=-0.5

because of x2<x1 the sequence monotonously decreases

Since for all n,xn>0, "\lim _{n\to \infty }\left(1+\frac{1}{n}\right)=1" then it is bounded below

{−1−1/ n ∨ n∈N}  

x1=-1-1=-2

x2=-1-1/2=-1.5

-1.5-(-2)=0.5

because of x2>x1 the sequence monotonously increases

Since for all n,xn>0 "\lim _{n\to \infty }\left(-1-\frac{1}{n}\right)=-1" then it is bounded above

which means that the limit of the set exists.